OpenStudy (anonymous):

What about this dy/dx= 1+2x^2/x cos y

OpenStudy (anonymous):

Is it like this?\[\frac{dy}{dx} = 1 + 2 \frac{x^2}{x \cos y}\]

OpenStudy (anonymous):

no, it's like this: dy/dx= (1+2x^2)/(x cos y)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{1+2x^2}{x \cos y}\] \[ \cos y dy = \frac{1}{x}dx + 2x dx\]

OpenStudy (anonymous):

Integrate it now! :-D

OpenStudy (anonymous):

I am not a Greedy kid (Or, at least I believe so) but if you'd be kind enough to award me a medal for my work (at least for the LaTeX), I will definitely level up soon.

OpenStudy (anonymous):

Hahah, I will but how ?

OpenStudy (anonymous):

Oh, I see. Just click on Good Answer